OpenStudy (anonymous):

help? @bibby @Data_LG2 @Directrix

OpenStudy (anonymous):

oh hold on let me get screenshot

OpenStudy (anonymous):

http://prntscr.com/6fplth

OpenStudy (anonymous):

it part a basically asking why the solution is where the lines intersect?

OpenStudy (bibby):

pretty much, yeah

OpenStudy (anonymous):

ok the wording is weird

OpenStudy (anonymous):

let me put together an explanation and can you tell me if it looks ok?

OpenStudy (bibby):

sure

OpenStudy (anonymous):

ok here The x-coordinates of the points to the solution to where the equations intersect is the solution of x, because that is where the two lines meet. i felt like i explained that weird. its awkward fr me to explain

OpenStudy (anonymous):

how would you explain it?

OpenStudy (anonymous):

because when you graph the two equations, you will find a coordinate at which they meet. This coordinate would have an x-coordinate and a y-coordinate which is your solution

OpenStudy (bibby):

I've never really had to explain it until I read that question. as I understand it, when you set 2 equations equal to each other, you're finding the x coordinate where they have an equal y for example, y=2x+6, y=6x-2 2x+6=6x-2 2x+8=6x 8=4x x=2

OpenStudy (bibby):

where they share a y*

OpenStudy (bibby):

so when we solve that equation, we've solved the x value of the intersection point

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

i got part b and c done

OpenStudy (bibby):

do you want me to give you one last run through on your answer for part a

OpenStudy (anonymous):

sure

OpenStudy (bibby):

did you rewrite it at all?

OpenStudy (anonymous):

yes sorry for late response